OpenStudy (anonymous):

Can the Double-Angle formula be used to find sin 22.5°? If so, show it.

OpenStudy (anonymous):

@animal_lover36 @BATMAN31 @cupcake111 @gorv @iGreen please? Help?

OpenStudy (batman31):

Taking a timed test, i can help if you'll wait an hour.

OpenStudy (gorv):

u know double angle formula

OpenStudy (gorv):

sin2x=sin22.5

OpenStudy (anonymous):

It's not a test. It's an assignment but I'd like it done as soon as possible. I have another class to do. The formula I think I have to use is sin2x = 2sin x cos x?

OpenStudy (gorv):

2x=22.5 x=22.5/2=11.25

OpenStudy (gorv):

but we dont know the value of 11.25

OpenStudy (anonymous):

So that's as far as it can go?

OpenStudy (gorv):

sin2x=2sinx*cosx=2*sin11.25*cos11.25

OpenStudy (anonymous):

Is that solvable?

OpenStudy (gorv):

well it will be said solvalble if we use standard value ...so i say no

OpenStudy (anonymous):

Okay well I appreciate it. :) I'll give you a medal for the help.